OpenStudy (diamond77108):

plz help me i dont know how to do this 4y-3=5y+2

alones (alones):

Alright so let's start from subtracting \( 5y\) from each side \[\ 4y−3−5y=5y+2−5y\] \[\ −y−3=2\]

alones (alones):

So than add \(\ 3\) to both sides \[\ −y−3+3=2+3\] \[\ −y=5\]

alones (alones):

At your last stepdivide by \(-1\) \[\frac{ -y }{ -1 }=\frac{ 5 }{ -1 }\]

OpenStudy (diamond77108):

Thx for showing me how to do it

alones (alones):

You're welcome and also your final answer would be \( y=−5\)